OpenStudy (anonymous):

Find the type and number of solutions for the given equation. 2x² + 5 = 2x

OpenStudy (shamil98):

2x^2 -2x + 5 = 0 b^2 - 4ac = discriminant

OpenStudy (shamil98):

the discriminant determines what type of solutions you will have

OpenStudy (anonymous):

-36?

OpenStudy (shamil98):

In this case 4 - 4(5)(2) = -36 A negative discriminant means no real solutions, and two complex solutions.

OpenStudy (anonymous):

so what would the type of number be?

OpenStudy (shamil98):

complex ..

OpenStudy (shamil98):

the numbers are complex as in with i

OpenStudy (anonymous):

so what would i put for an answer? Like, i understand how to do the work, but what to put for the answer is where i'm bunked

OpenStudy (shamil98):

imaginary numbers, two complex solutions.

OpenStudy (anonymous):

\[2x^2 - 3x = 8\] \[2x^2 -3x -8 = 0\] \[b^2 - 4ac = (-3)^2 - 4(2)(-8)\] \[9 + 64 = 73\] @shamil98

OpenStudy (shamil98):

If the discriminant is greater than zero that are two real solutions.

OpenStudy (anonymous):

so what would i write for the answer? "two real solutions"?

OpenStudy (shamil98):

Yes.

OpenStudy (anonymous):

mind helpin me out with a couple more? @shamil98

OpenStudy (shamil98):

ask them in the question box and close this one ,please.

OpenStudy (anonymous):

okay, MEDALS AWAY!!!!!